OpenStudy (mew55):

The Web page for a public library had a mean of 257 hits per day with a standard deviation of 37. What percentage of days had less than 187 hits?

OpenStudy (kirbykirby):

I am guessing it is assumed to follow a normal distribution?

OpenStudy (kirbykirby):

Well assuming it is, We want \(P(X<187)\). We can standardize this: \[P\left(\frac{X-257}{37}<\frac{187-257}{37}\right)=P(Z<-1.89)\]

OpenStudy (kirbykirby):

Do you know what to do from here?